OpenStudy (anonymous):

solve system of equations using substitution: 5x+6y+8=0 3x-2y+16=0 Please show work (:

OpenStudy (anonymous):

er why substitution. looks like subtraction would be wayyy easier.

OpenStudy (anonymous):

substitue substituion with subtraction :)

OpenStudy (anonymous):

substitute*

OpenStudy (anonymous):

multiply the second equation by 3 and subtract it from the first then solve for x (that's how id do it)

OpenStudy (anonymous):

I have to do substitution. Actually, I got an answer to this but it didn't look right. So thats why I'm asking about it.

OpenStudy (anonymous):

so you get 9x+6y+48 and subtracting this from 5x+6y+8 = 0 and you get -4x-40=0 so you get x =10

OpenStudy (anonymous):

sorry that first equation should say =0

OpenStudy (anonymous):

but it's got to be substitution. I have to show my work.

OpenStudy (anonymous):

kk. substitute 6y (solve from 3 times the second equation) into the first equation

OpenStudy (anonymous):

6y=48-9x

OpenStudy (anonymous):

where did you get the 48 and nine from??

OpenStudy (anonymous):

i multiplied the whole equation by three

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

it makes it easier so you dont have to divide anything it's also legal because i did it to both sides of the equation

OpenStudy (anonymous):

Alright, but you still have to divide tot get y alone

OpenStudy (anonymous):

you can but it's easier and still technically substitution. I copied the problem down work 6y =48+9x

OpenStudy (anonymous):

but wouldn't something be negative?

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

nope the negatives cancel out on the second equation when you solve for y

OpenStudy (anonymous):

or 6y

OpenStudy (anonymous):

alright, so i got y=9/6x+8

OpenStudy (anonymous):

so you end up with 5x + 48+9x+8=0 after subbing in 6y

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

that looks right

OpenStudy (anonymous):

how? if you use what i did to substitute for y then there can't be 48 or nine in it.

OpenStudy (anonymous):

i did it differently than you, instead of solving for y i solved for 6y and substitute what 6y is equal to in terms of x according the the second equation into the first equation

OpenStudy (anonymous):

either way is valid

OpenStudy (anonymous):

you need to do it schuyler's way if you want to work with whole numbers. else you get nasty fractions. Which is why you should use subtraction in these instances.

OpenStudy (anonymous):

Yeah but I can't I have to show work for subsitition. and I don't understand schuyler's way at all

OpenStudy (anonymous):

ight check this out yo, homz.

OpenStudy (anonymous):

it's okay if you don't understand yet. Just think about it though. You don't necessarily have to solve for just the variable. you can solve for whatever you like. 6y for example instead of y. as long as you follow the rules of algebra you'll get a valid answer, and with enough practice you'll get it

OpenStudy (anonymous):

if you multiply the second equation by 3. Everything gets trippled. so instead of 3x-2y+16, you get 9x-6y+48

OpenStudy (anonymous):

alrighttt.... i guess I kind of get it. so you replace 6y with what the other 6y equals?

OpenStudy (anonymous):

so now your system looksalikedis. 5x+6y+8=0 9x-6y+48= 0

OpenStudy (anonymous):

now when your ststem lookalikedat, you can add everything together :O (so this means you'll get 14x +nothing because +6+-6=0 +54 because 48+8 = 54 leaving you with 14x+54=0

OpenStudy (anonymous):

with 14x+54 = 0 you subtract the 54 over to the left, leaving you with 14x = -54, then divide -54 by 14x, leaving you with - 4. So X now = -4!!!!! now we can substitute.

OpenStudy (anonymous):

now I'm even more confused. Then I was before. Thanks for trying though.

OpenStudy (anonymous):

aw.

OpenStudy (anonymous):

okay! I shall devise a plan.

OpenStudy (anonymous):

alright. kind of sounds a bit devious... (:

OpenStudy (anonymous):

what's 56 divided by 14?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

but it should be -4 when you subtract 54 to the right side

OpenStudy (anonymous):

alright I think I got it! is x= -4 !

OpenStudy (anonymous):

YES.

OpenStudy (anonymous):

I drew you a big picture :) want it or do you understand?

OpenStudy (anonymous):

but now i don't know how to do the next part

OpenStudy (anonymous):

HOKAY. Just wait :P

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

oh crap I can't send it okay I'll explain it

OpenStudy (anonymous):

you know we got x=-4 right?

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

so going back to the equations. You started with 5x+6y+8 = 0 right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

:o but you know what x is!! so it's really 5 TIMES -4 +6y+8 = 0

OpenStudy (anonymous):

:O!!!!!!!!!!!!!!

OpenStudy (anonymous):

I GOT IT! I think! x=-4 and y=14

OpenStudy (anonymous):

?

OpenStudy (anonymous):

well 5 x -4 is -20. -20 +8 is -12 :P

OpenStudy (anonymous):

so if you have 6y-12 = 0 what do you do?

OpenStudy (anonymous):

wait why are you using that problem? shouldn't you be using 6y=9x+48?

OpenStudy (anonymous):

noooooo. You should be plugging it back into the ORIGINAL equation

OpenStudy (anonymous):

HOWEVER

OpenStudy (anonymous):

Oooooohhh. weellll theeen never mind!

OpenStudy (anonymous):

9(-4) = -36 +48 = 12 all the same ;P

OpenStudy (anonymous):

so you have 6y-12 = 0. we want to add 12 to both sides of the =

OpenStudy (anonymous):

wait so is y=2

OpenStudy (anonymous):

yes y is 2 but i meant to say you are subtracting 12 from both sides

OpenStudy (anonymous):

negative 2!!!

OpenStudy (anonymous):

its fine! I totally got it now!

OpenStudy (anonymous):

THanks

OpenStudy (anonymous):

it's -2!!! lol sorry

OpenStudy (anonymous):

glad I could help!

OpenStudy (anonymous):

wait! Why is it negative if you add 12?

OpenStudy (anonymous):

because I fluttere dup. you're supposed to subtract 12

OpenStudy (anonymous):

why its a negative 12?

OpenStudy (anonymous):

6y+12 = 0 , so you -12 to get 6y= -12. then -12/6 = -2

OpenStudy (anonymous):

no because -4 x 5 = -20 + 8 = -12

OpenStudy (anonymous):

Multiply the first equation by 2. You now have two equations where the constant is 16. Equate the x and y terms in the first equation to the x and y terms in the second equation and solve for y.\[10x+12y\text{=}3x-2y \]\[y=-\frac{x}{2} \]Plug y into the second equation and solve for x.